In the CIA World Fact Book, GDP per capita in the United States in 2010 was approximately $47,400. The formula for growth used in that spreadsheet for any given year, yt, is yt = y0 (1 + g)t, where y0 is the value of GDP in the beginning year, yt is the value of GDP for the specific year in question, and t is the number of years after y0. If y0 is GDP per capita in 2010 and the economy continues to grow at approximately 3% as it did in 2010, what will be the value of GDP per capita in ten years?


1
Expert's answer
2021-10-11T10:53:17-0400

Value of GDP per capita in 10 years i.e value of GDP per capita in 2020 is given by,

"y_{10} = y0 (1+gy)^{10}" ( t=10 since y0 is the GDP per capita in 2010)

"y_{10} = \$47,400(1+0.03)^{10}\\

y_{10} = \$ 47,400\times 1.3439\\

= \$ 63,700.86"

Therefore, GDP per capita in 2020"= \$ 63,701"